Brief Exercise 11-3 Oriole Company purchased machinery on January 1, 2017, for $90,400. The machinery is estimated to have aBrief Exercise 11-3 Oriole Company purchased machinery on January 1, 2017, for $90,400. The machinery is estimated to have a salvage value of $9,040 after a useful life of 8 years. Compute 2017 depreciation expense using the sum-of-the-years'-digits method. Depreciation expense $ LINK TO TEXT Compute 2017 depreciation expense using the sum-of-the-years'-digits method, assuming the machinery was purchased on April 1, 2017. (Round answer to 0 decimal places, e.g. 5,125.) Depreciation expense $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

1) Sum of year digit = 8+7+6+5+4+3+2+1 = 36

Depreciation expense = (90400-9040)*8/36 = 18080

2) Depreciation expense 2017 = (90400-9040)*8/36*9/12 = 13560